# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v2

@@ 3.1 Linear Least-squares with Gaussian Data
independently for each index $i$, where $Q\in\mathbb{R}^{(n+1)\times(n+1)}$ is a positive semidefinite matrix. Furthermore, let $Q_{n}$ be the top left $n\times n$ block of $Q$, assume for simplicity that $Q_{n}$ is full rank, and let $Q_{n}=LL^{\top}$ be its Cholesky decomposition. Denote the singular values of $L$ by $\sigma_{1}\geq\dots\geq\sigma_{n}>0$. Finally, denote by $y$ the solution to the underlying statistical problem, which is unique since $Q_{n}$ is full rank:
(17) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$
@@
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the randomly generated finite problem (16), where $m$ denotes the total number of generated samples and $k$ is the block size. Then the results of Theorem 3.1 hold with
(18) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
In addition, the convergence parameter $\alpha$ satisfies
(19) $\lim_{m\rightarrow\infty}\alpha\geq C_{n,k}\max\left\{\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}},\max_{2\leq\ell<k}\frac{(\ell-1)\sigma_{n}^{2}}{\sum_{i\geq k-\ell-1}\sigma_{i}^{2}}\right\}$
with $C_{n,k}\rightarrow 1$ as $n\rightarrow\infty$ for fixed $k$.
Furthermore, as long as $k\geq 6$ and $\operatorname{rank}(L)\geq 2k$, the variance term satisfies
(20) $\lim_{m\rightarrow\infty}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}\leq\frac{200}{\sigma_{2k}^{2}}\cdot\lim_{m\rightarrow\infty}\frac{\left\lVert r\right\rVert^{2}}{m}.$
In the above theorem, the randomness in the Gaussian data is handled by first assuming a fixed number of samples $m$, and then considering the asymptotic limit $m\rightarrow\infty$. The result is that in the case of Gaussian data, the ordinary least-squares solution is recovered, the convergence rate $\alpha$ improves at least linearly with the block size $k$, and the variance of the iterates is bounded. In addition, the following corollary, which is based on Corollary 3.4 of [14], shows that RBK converges much faster than mSGD for polynomially decaying singular values.
@@
Consider the setting of Theorem 3.2 with fixed $k\leq n/2$, and assume the $L$ factor has polynomial spectral decay $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$ and some $\beta>1$. Then the convergence parameters of RBK and mSGD satisfy
(21) $\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq Ck^{\beta}\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}},\;\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq k\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}}$
@@
Similarly, the faster convergence rate of RBK over mSGD extends to exponentially decaying singular values. The proofs of Theorems 3.2 and 3.3, provided in Appendix B, rely on a connection between RBK-U for Gaussian data and sketch-and-project with Gaussian sketch matrices. These results generalize the techniques of [13, 14] to the case of inconsistent linear systems, improving upon the results of [37] in terms of both the convergence rate and the variance. It is worth noting that Gaussian data is just one way to generate a benign or “incoherent” problem. We expect that Theorems 3.2 and 3.3 could be generalized to other cases such as problems that have been preprocessed using a randomized Hadamard transform, potentially extending the results of [15] to the inconsistent case.
3.2 Implementation Details
To stably implement the RBK iteration (5), we employ an SVD-based least-squares solver to calculate $A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$ The most expensive part of this procedure is the SVD, which has an asymptotic cost of $O(nk^{2})$. This cost is greater than the cost of an mSGD iteration, which is $O(nk)$. Whether it is worthwhile to expend this extra cost depends on the structure of the matrix $A$, and especially on its singular values as discussed in the previous section.
@@ 3.3 Numerical Demonstration
We verify our theoretical results for the RBK-U and mSGD algorithms on two problems with Gaussian data, with results in Fig. 1. We apply tail averaging to each algorithm to observe convergence beyond the variance horizon. As expected, tail-averaged RBK-U (TA-RBK-U) converges much more rapidly than tail-averaged mSGD (TA-mSGD) in the presence of fast singular value decay. We additionally explore the effect of choosing different burn-in periods for TA-RBK-U, with results in Fig. 2. As expected, optimal performance is attained by setting $T_{b}$ to be just slightly after the iterates reach their fixed convergence horizon. The code for all of our experiments is linked in the supplementary material, and further details on these particular experiments can found in F.1.
Figure 1: Comparison of methods on two problems with Gaussian data, with no singular value decay (left) and rapid singular value decay (right). The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
Figure 2: Behavior of TA-RBK-U with different burn-in periods $T_{b}$. The problem and the reported error quantity are the same as in the left panel of Fig. 1.
@@ 4 Robustness through Regularization
To address the shortcomings of RBK-U in the case of general data, we propose to incorporate regularization into the RBK algorithm. A natural way to do this is to replace the RBK iteration with a stochastic proximal point iteration [2], namely
@@
(25) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\frac{\left\lVert r^{(\rho)}\right\rVert^{2}}{m}.$
Finally, when the squared row norms of $A$ are uniformly bounded by a constant $N$, then $\kappa(\overline{W})\leq 1+\frac{N}{\lambda}$ and the weighted residual $r^{(\rho)}$ can be bounded in terms of the ordinary least-squares residual $r$:
@@
(26) $\left\lVert r^{(\rho)}\right\rVert^{2}\leq\left(1+\frac{N}{\lambda}\right)\left\lVert r\right\rVert^{2}.$
@@
Note that the values of $x^{(\rho)}$ and $\alpha$ here differ from those in Section 2 due to the different choice of $M(A_{S})$ used by ReBlocK. The advantage relative to RBK-U is that ReBlocK-U is able to control both the variance of the iterates and the conditioning of the weight matrix $\overline{W}$ in terms of the reciprocal of the regularization parameter $\lambda$. As a result, the algorithm converges robustly even when the problem contains many nearly singular blocks $A_{S}$. It is worth noting that these advantages could also be attained by truncating the singular values that are smaller than $\lambda$ in the RBK iteration (5). However, the ReBlocK iteration is additionally justified by the fact that it is cheaper to implement than the RBK iteration; see Section 4.1 and Fig. 8 for further discussion of this point.
@@
The proof of Theorem 4.1 is provided in Appendix C. Relative to the proof of Theorem 3.1, the proof of Theorem 4.1 is more complicated because the ReBlocK iteration does not lead to an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$. Instead, the proof relies on a bias-variance decomposition inspired by [12, 23, 18], which also leads to the extra factors of $2$ in the convergence bound.
@@
We are not yet able to analyze the convergence rate parameter $\alpha$ of ReBlocK-U, even in the case of Gaussian data, as to our knowledge there is no existing work bounding the quality of a regularized Gaussian sketch. However, convergence to the ordinary least-squares solution is obtained for a broader class of noisy linear least-squares problems; see Appendix D. Additionally, a fast rate of convergence to the ordinary least-squares solution can be shown when sampling from an appropriate determinantal point process; see Appendix E.
@@ 4.1 Implementation Details
To implement the ReBlocK iterations (7), we directly calculate the $k\times k$ matrix $A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI$ and then use a Cholesky-based linear system solver to calculate $(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x)$. This computation is stable as long as $\lambda$ is not chosen to be too small. Using this approach, the most expensive part of the ReBlocK iteration is calculating $A_{S_{t}}A_{S_{t}}^{\top}$, which has an asymptotic cost of $O(nk^{2})$ just like RBK. Nonetheless, in practice the matrix-matrix multiplication for ReBlocK can be significantly cheaper than the SVD used for RBK. For example, in the experiments of Section 5, ReBlocK iterations are over twenty-five times faster than RBK iterations, as reported in Fig. 8. For the largest problems, the ReBlocK iterations could be further accelerated using iterative solvers; see for example Section 4.1 of [16].
4.2 Numerical Demonstration
We now test the ReBlocK-U algorithm on two problems whose columns are discretized representations of continuous functions, leading the matrices $A$ to contain many nearly singular blocks of rows. We set $\lambda=0.001$ and generate the inconsistency using random Gaussian noise. We find that TA-RBK-U is unstable for these problems, as expected. Furthermore, TA-ReBlocK-U converges much faster than TA-mSGD in the presence of rapid singular value decay. To understand the impact of the inconsistency on the algorithms, we re-run the case with rapid singular value decay with varying levels of inconsistency, with results in Fig. 4. We find qualitatively similar results in every case. Further details on these experiments can be found in F.1.
Figure 3: Comparison of methods on two inconsistent problems with many nearly singular blocks, with mild singular value decay (left) and rapid singular value decay (right). The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
Figure 4: Exploration of the effect of the inconsistency on the behavior of the algorithms. The variance of each entry of the noise vector is set to $61e-$ 6 (top left), $41e-$ 4 (top right), and $21e-$ 2 (bottom). The top right panel here is identical to the right panel of Fig. 3, and is reproduced here only for comparison with the top left and bottom.
@@ 5 Natural Gradient Optimization
Our original motivation for this work comes from the problem of training deep neural networks using natural gradient descent [1], which is based on an efficient natural gradient induced by a problem-dependent Riemannian metric. Natural gradient descent has been studied extensively in the machine learning community; see [28, 38, 27]. Furthermore, there is increasing evidence that natural gradient methods can improve the accuracy when training neural networks to solve physical equations. See [29, 9] for applications to physics-informed neural networks and [35, 40] for applications to neural network wavefunctions.
@@
Calculating the natural gradient direction $G_{N}$ using (28) requires a linear solve against the $n\times n$ matrix $F$, which is very challenging in realistic settings when $n\geq 10^{6}.$ This has motivated the development of approximate schemes such as [28]. An alternative approach is to reformulate $G_{N}$ using the structure of $F$ and $\nabla_{\theta}L(\theta)$:
@@
where the norm in the final expression is the $L_{2}$-norm in function space. This least-squares formulation has been pointed out for example by [27, 7, 19], with the work of Chen and Heyl empowering major advances in the field of neural quantum states. A major goal of the current work is to provide a more solid foundation for the development of natural gradient approximations along these lines.
The natural way to access the data when training a neural network is to sample a set of points in the domain $\Omega$ and evaluate the target function $f$, the network outputs $f_{\theta}$, and the network gradients $\nabla_{\theta}f_{\theta}$ at the sampled points. This is precisely equivalent to sampling a small subset of the rows of (32), which motivates the consideration of row-access least-squares solvers for calculating natural gradient directions. Furthermore, both empirical evidence from scientific applications [33, 45] and theoretical evidence from the literature on neural tangent kernels [4, 39, 5] suggest that $J$ should be expected to exhibit fast singular value decay, motivating the possibility that RBK and ReBlocK could converge rapidly when solving (32). Note that these observations translate straightforwardly from the simple function learning problem (27) to the realistic problems of training physics-informed neural neteworks or neural network wavefunctions.

### v3

@@ 4 Linear Least-squares with Gaussian Data
where $Q\in\mathbb{R}^{(n+1)\times(n+1)}$ is a positive semidefinite covariance matrix. Furthermore, let $Q_{n}$ be the top left $n\times n$ block of $Q$, assume for simplicity that $Q_{n}$ is full rank, and let $Q_{n}=L_{n}L_{n}^{\top}$ be its Cholesky decomposition. Denote the singular values of $L_{n}$ by $\sigma_{1}\geq\dots\geq\sigma_{n}>0$.
@@
Consider the equivalent of the RBK-U algorithm for the statistical least-squares problem Eq. 19, for which each sample $A_{S_{t}},b_{S_{t}}$ is drawn by generating $k$ independent samples from $\mathcal{N}(0,Q)$. Then the results of Theorem 3.1 apply and it holds $x^{(\rho)}=x^{*}$. In addition, the convergence parameter $\alpha$ satisfies
(20) $\alpha\geq C_{n,k}\max\left\{\frac{k\sigma_{n}^{2}}{\|L_{n}\|_{F}^{2}},\max_{2\leq\ell<k}\frac{(\ell-1)\sigma_{n}^{2}}{\sum_{i\geq k-\ell-1}\sigma_{i}^{2}}\right\}$
with $C_{n,k}\rightarrow 1$ as $n\rightarrow\infty$ for fixed $k$. Furthermore, as long as $k\geq 6$ and $\operatorname{rank}(L_{n})\geq 2k$, the variance term satisfies
(21) $V\leq\frac{200}{\sigma_{2k}^{2}}\cdot\mathbb{E}_{[a^{\top}\;b]\sim\mathcal{N}(0,Q)}\left[(a^{\top}x^{*}-b)^{2}\right].$
In summary, in the case of Gaussian data the ordinary least-squares solution is recovered, the convergence rate $\alpha$ improves at least linearly with the block size $k$, and the variance of the iterates is bounded. In addition, the following corollary, which is based on Corollary 3.4 of [derezinski2024sharp], shows that RBK converges much faster than mSGD for polynomially decaying singular values.
@@
Consider the setting of Theorem 4.1 with fixed $k\leq n/2$, and assume the $L_{n}$ factor has polynomial spectral decay $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$ and some $\beta>1$. Then the convergence parameters of RBK and mSGD satisfy
(22) $\alpha^{\rm RBK}\geq Ck^{\beta}\tfrac{\sigma_{n}^{2}}{\|L_{n}\|_{F}^{2}},\;\alpha^{\rm mSGD}\leq k\tfrac{\sigma_{n}^{2}}{\|L_{n}\|_{F}^{2}}$
@@
Similarly, the faster convergence rate of RBK over mSGD extends to exponentially decaying singular values. The proofs of Theorems 4.1 and 4.2, provided in Appendix B, rely on a connection between RBK-U for Gaussian data and block Gaussian Kaczmarz for arbitrary data. These results generalize the techniques of [derezinski2024sharp] to the case of inconsistent linear systems, improving upon the results of [rebrova2021block] in terms of both the convergence rate and the variance.
@@ 4.1 Numerical Demonstration
We verify our theoretical results for the RBK-U and mSGD algorithms on two problems with Gaussian data, with results in Fig. 3. We apply tail averaging to each algorithm to observe convergence beyond the variance horizon. As expected, tail-averaged RBK-U (TA-RBK-U) converges much more rapidly than tail-averaged mSGD (TA-mSGD) in the presence of fast singular value decay. We additionally explore the effect of choosing different burn-in periods for TA-RBK-U, with results in Fig. 4. As expected, optimal performance is attained by setting $T_{b}$ to be just slightly after the iterates reach their fixed convergence horizon. More details on these experiments can be found in Appendix F.
Figure 3: Comparison of methods on two problems with Gaussian data, with no singular value decay (left) and rapid singular value decay (right). The vertical dotted line indicates the burn-in time, before which results are shown for individual iterates.
Figure 4: Behavior of TA-RBK-U with different burn-in periods $T_{b}$. The problem is the same as in the left panel of Fig. 3.
@@ 5 Robustness through Regularization
To address the shortcomings of RBK-U in the case of general data, we propose to incorporate regularization into the RBK algorithm. A natural way to do this is to replace the RBK iteration with a stochastic proximal point iteration [asi2019stochastic], namely
@@
(26) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{4}{\alpha^{2}(T-T_{b})}V.$
with $V=\mathbb{E}\,_{S\sim\rho}\|A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}\|^{2}$. Finally, the condition number $\kappa(\overline{W})$, the weighted residual $r^{(\rho)}$, the bias $x^{(\rho)}-x^{*}$, and the variance $V$ are bounded as
(27) $\kappa(\overline{W})$ $\leq 1+\frac{1}{\lambda}\cdot\max_{i}\|a_{i}\|^{2},$
@@
$\|r^{(\rho)}\|$ $\leq\sqrt{\kappa(\overline{W})}\cdot\left\lVert r\right\rVert,$
$\|x^{(\rho)}-x^{*}\|$ $\leq\sqrt{\kappa(\overline{W})-1}\cdot\left\lVert A^{+}\right\rVert\cdot\|r\|\leq\frac{1}{\sqrt{\lambda}}\cdot\kappa(A)\cdot\|r\|,$
$V$ $\leq\frac{1}{4\lambda}\cdot\kappa(\overline{W})\cdot\frac{\left\lVert r\right\rVert^{2}}{m}.$
@@
Note that the values of $x^{(\rho)}$, $\alpha$, and $V$ here differ from those in Section 2 due to the different choice of $M(A_{S})$ used by ReBlocK. The advantage relative to RBK-U is that ReBlocK-U is able to control the condition number of the weight matrix, the weighted residual, the bias of the weighted least-squares solution, and the variance of the iterates in terms of the reciprocal of the regularization parameter $\lambda$. As a result, the algorithm converges robustly even when the problem contains many nearly singular blocks $A_{S}$.
As a concrete example of the benefits of regularization, recall the isosceles triangle problem Eq. 17 for which the bias and variance of RBK-U both diverge as $\epsilon\rightarrow 0$. In contrast, for ReBlocK-U we have the following corollary:
Corollary 5.2.
Consider the isosceles triangle problem Eq. 17 from Equation 18. For ReBlocK-U with any fixed $\lambda>0$, the bounds Eq. 27 imply
(28) $\lim_{\epsilon\rightarrow 0^{+}}\|r^{(\rho)}\|=\lim_{\epsilon\rightarrow 0^{+}}\|x^{(\rho)}-x^{*}\|=\lim_{\epsilon\rightarrow 0^{+}}V=0$
since the row norms $a_{i}$ remain bounded and the residual norm $\left\lVert r\right\rVert$ vanishes as $\epsilon\rightarrow 0$.
In other words, introducing even a small amount of regularization prevents the algorithm from failing catastrophically. We confirm this effect numerically by directly calculating the bias for various values of $\epsilon$ and $\lambda$, with results in Figure 5. Note that for fixed $\epsilon$, the bias decreases monotonically with $\lambda$, as expected.
Figure 5: Size of the bias $x^{(\rho)}-x^{*}$ for ReBlocK-U with $k=2$ on the isosceles triangle problem of Eq. 18, for various values of $\epsilon$ and $\lambda$. Left: for $\lambda=0$ the bias grows without bound as $1/\epsilon\rightarrow\infty$, whereas for any $\lambda>0$ the bias reaches a fixed maximum value and then decays to zero. Right: for any fixed $\epsilon$ the bias decreases monotonically with $\lambda$.
@@
It is worth noting that these advantages could also be attained by truncating the small singular values in the RBK iteration (5). However, the ReBlocK iteration is additionally justified by the fact that it is cheaper to implement than the RBK iteration; see Section 5.2 and Fig. 12 for further discussion of this point. The proof of Theorem 5.1 is provided in Appendix C. Relative to the proof of Theorem 3.1, the proof of Theorem 5.1 is more complicated because the ReBlocK iteration does not lead to an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$. Instead, the proof relies on a bias-variance decomposition inspired by [defossez2015averaged, jain2018parallelizing, epperly2024randomizedkaczmarztailaveraging], which also leads to the extra factors of 22 in the convergence bound.
@@
We are not yet able to analyze the convergence rate parameter $\alpha$ of ReBlocK-U, even in the case of Gaussian data, as to our knowledge there is no existing work bounding the quality of a regularized Gaussian sketch. However, a fast rate of convergence to the ordinary least-squares solution can be shown when sampling from an appropriate determinantal point process; see Appendix E. Additionally, convergence to the ordinary least-squares solution is obtained for a broader class of noisy linear least-squares problems; see Appendix D.
5.1 Numerical Demonstration
We now demonstrate that the benefits of regularization can translate to more realistic types of problems. To do so we revisit the examples from Figure 2, for which RBK-U exhibited severe instabilities. We observe in Figure 6 that ReBlocK-U with $\lambda=1e-3$ is stable for both of these problems and converges much faster than TA-mSGD in the presence of rapid singular value decay. To understand the impact of the inconsistency on the algorithms, we re-run the case with rapid singular value decay with varying levels of inconsistency, with results in Figure 7. We find qualitatively similar results in every case. Further details on these experiments can be found in Appendix F.
Figure 6: Performance of ReBlocK-U for the same two inconsistent problems from Figure 2. These problems both contain many nearly singular blocks, with mild singular value decay (left) and rapid singular value decay (right). The vertical dotted line indicates the burn-in time, before which results are shown for individual iterates.
Figure 7: Exploration of the effect of the inconsistency on the behavior of the algorithms. The variance of each entry of the noise vector is set to 1e61e-6 (top left), 1e41e-4 (top right), and 1e21e-2 (bottom). The top right panel here is identical to the right panel of Figure 6, and is reproduced here only for comparison with the top left and bottom
@@ 5.2 Implementation Details
To implement the ReBlocK iterations (7), we directly calculate the $k\times k$ matrix $A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI$ and then use a Cholesky-based linear system solver to calculate $(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x)$. This computation is stable as long as $\lambda$ is not chosen to be too small. Using this approach, the most expensive part of the ReBlocK iteration is calculating $A_{S_{t}}A_{S_{t}}^{\top}$, which has an asymptotic cost of $O(nk^{2})$ just like RBK. Nonetheless, in practice the matrix-matrix multiplication for ReBlocK can have a much smaller preconstant than the QR decomposition used for RBK. For example, in the experiments of Section 6, ReBlocK iterations are over twenty-five times faster than RBK iterations, as reported in Fig. 12. We emphasize that this gap is not related to the use of the Cholesky decomposition but instead arises because ReBlocK avoids doing any matrix decomposition at all directly on the $k\times n$ matrix $A_{S}$. For the largest problems, the ReBlocK iterations could be further accelerated using iterative solvers; see for example Section 4.1 of [derezinski2025].
@@ 6 Natural Gradient Optimization
Our original motivation for this work comes from the problem of training deep neural networks using natural gradient descent [amari1998natural], which is based on an efficient natural gradient induced by a problem-dependent Riemannian metric. Natural gradient descent has been studied extensively in the machine learning community; see [martens2015optimizing, ren2019efficient, martens2020new]. Furthermore, there is increasing evidence that natural gradient methods can improve the accuracy when training neural networks to solve physical equations. See [muller2023achieving, dangel2024kronecker] for applications to physics-informed neural networks and [pfau2020ab, schatzle2023deepqmc] for applications to neural network wavefunctions.
@@
Calculating the natural gradient direction $G_{N}$ using (30) requires a linear solve against the $n\times n$ matrix $F$, which is very challenging in realistic settings when $n\geq 10^{6}.$ This has motivated the development of approximate schemes such as [martens2015optimizing]. An alternative approach is to reformulate $G_{N}$ using the structure of $F$ and $\nabla_{\theta}L(\theta)$:
@@
where the norm in the final expression is the $L_{2}$-norm in function space. This least-squares formulation has been pointed out for example by [martens2020new, chen2024empowering, goldshlager2024kaczmarz], with the work of Chen and Heyl empowering major advances in the field of neural quantum states. A major goal of the current work is to provide a more solid foundation for the development of natural gradient approximations along these lines.
The natural way to access the data when training a neural network is to sample a set of points in the domain $\Omega$ and evaluate the target function $f$, the network outputs $f_{\theta}$, and the network gradients $\nabla_{\theta}f_{\theta}$ at the sampled points. This is precisely equivalent to sampling a small subset of the rows of (34), which motivates the consideration of row-access least-squares solvers for calculating natural gradient directions. Furthermore, both empirical evidence from scientific applications [park2020geometry, wang2022and] and theoretical evidence from the literature on neural tangent kernels [bietti2019inductive, ronen2019convergence, cao2019towards] suggest that $J$ should be expected to exhibit fast singular value decay, motivating the possibility that RBK and ReBlocK could converge rapidly when solving (34). Note that these observations translate straightforwardly from the simple function learning problem (29) to the realistic problems of training physics-informed neural networks or neural network wavefunctions.
It is also worth noting that while we focus on the problem of calculating natural gradient directions, similar linear systems also arise when using neural networks to simulate the time evolution of either differential equations [bruna2024neural] or quantum systems [schmitt2020quantum]. When simulating time dynamics it is essential to solve each linear system to high accuracy. This provides additional motivation for developing block row access methods that can solve such linear systems accurately without preprocessing.